In the nucleus of an atom, two protons are separated by a distance of 1 x 10 -5m. What is the magnitude of the electric force between them?
Answer #1

Electrostatic force is given by:

F = k*q1*q2/r^2

k = 9*10^9

q1 = q2 = charge on proton = +1.6*10^-19 C

r = distance between two protons = 1*10^-5 m

So Using these values:

F = 9*10^9*(1.6*10^-19)^2/(1*10^-5)^2

F = 2.30*10^-18 N

Since both charge are positive, So force between them will be repulsive.

Please Upvote.